# Tele-LLMs: A Series of Specialized Large Language Models for Telecommunications

### v2

@@ 8.2. Pretraining Data Impact
Figure 7. Generated responses from different models for a telecommunications-themed prompt.
Figure 8. Training loss during instruction fine-tuning for the Tele-LLMs series.
@@
To illustrate this, let us consider two models: Gemma-2B and Gemma-2B-it [^9] https://huggingface.co/google/gemma-2b-it. The latter is an instruct version of the former that has been post-trained on instruction-following data, which includes tasks like question answering and summarization. As Table 5 shows, instruct models outperform their base counterparts on Tele-Eval. This is because the additional instruction-focused training allows the LLM to better handle questions, follow instructions, and leverage its internalized knowledge to effectively answer queries. This advantage is significant enough that Gemma-2B-it is able to outperform our telecommunications-adapted LLM, Gemma-2B-Tele. However, since Gemma-2B-Tele has been enriched with telecommunications knowledge, we can surpass the performance of Gemma-2B-it by applying a similar post-training instruction-following adaptation using a dataset like Alpaca [^10] https://huggingface.co/datasets/tatsu-lab/alpaca and utilizing the same training settings previously detailed, thus creating Gemma-2B-Tele-Alpaca.
@@
Table 5. LLM-Eval performance on Tele-Eval.
The significance of the above observation is that some base models have been trained on instruct-like data during their pretraining stage. An example of this is Microsoft Phi-1.5. In their paper (et al., 2023), the authors specify that Phi-1.5 can be prompted as an instruct model because the pretraining data included a large portion of question-answer formats. This leads to noticeable behavioral differences between Gemma-2B and Phi-1.5, as illustrated in Fig. 7. One can see that Phi-1.5 behaves more like a chatbot, following the instructions of the prompt, while Gemma-2B fills in sentences without treating the input prompt as an instruction to follow.
@@
Table 6. LLM-Eval performance on Tele-Eval.
8.3. Instructions Fine-tuning
@@
Training Settings. We maintain the training settings from Section 7 with minor adjustments. Specifically, we decrease the batch size to 128k tokens, set the context length to 2048, and limit the number of epochs to 1. For this stage, we use two datasets: Alpaca and Open-Instruct [^11] https://huggingface.co/datasets/VMware/open-instruct. The combined datasets provide approximately 200k samples for training.
Results. As seen in Fig. 8, the training loss consistently decreases as the model becomes more adept at following instructions. Next, in Table 6, we compare the Tele-Instruct versions of our models to their general instruct counterparts. Our instruction-tuned series outperforms the general instruct models, thereby showcasing their superior knowledge in telecom compared to their general-purpose counterparts.
9. Conclusions

### v3

@@ VIII-B Pretraining Data Impact
Figure 7: Generated responses from different models for a telecommunications-themed prompt.
@@
To illustrate this, let us consider two models: Gemma-2B and Gemma-2B-it [^9] https://huggingface.co/google/gemma-2b-it. The latter is an instruct version of the former that has been post-trained on instruction-following data, which includes tasks like question answering and summarization. As Table V shows, instruct models outperform their base counterparts on Tele-Eval. This is because the additional instruction-focused training allows the LLM to better handle questions, follow instructions, and leverage its internalized knowledge to effectively answer queries. This advantage is significant enough that Gemma-2B-it is able to outperform our telecommunications-adapted LLM, Gemma-2B-Tele. However, since Gemma-2B-Tele has been enriched with telecommunications knowledge, we can surpass the performance of Gemma-2B-it by applying a similar post-training instruction-following adaptation using a dataset like Alpaca [^10] https://huggingface.co/datasets/tatsu-lab/alpaca and utilizing the same training settings previously detailed, thus creating Gemma-2B-Tele-Alpaca.
@@
TABLE V: LLM-Eval performance on Tele-Eval.
The significance of the above observation is that some base models have been trained on instruct-like data during their pretraining stage. An example of this is Microsoft Phi-1.5. In their paper [2], the authors specify that Phi-1.5 can be prompted as an instruct model because the pretraining data included a large portion of question-answer formats. This leads to noticeable behavioral differences between Gemma-2B and Phi-1.5, as illustrated in Fig. 7. One can see that Phi-1.5 behaves more like a chatbot, following the instructions of the prompt, while Gemma-2B fills in sentences without treating the input prompt as an instruction to follow.
@@
TABLE VI: LLM-Eval performance on Tele-Eval.
VIII-C Instructions Fine-tuning
TABLE VII: Performance comparison of individually tuned models vs. unified instruction tuning
{NiceTabular}
c|c|c|cc|cc|cc|cc Models Citation Link Prediction Citation Recommendation Title Generation   Abstract Completion   Citation Sentence Generation   Introduction to Abstract
Accuracy Accuracy Precision F1 Score Precision F1 Score Precision F1 Score Precision F1 Score
Llama-3.2-1B 21.34 10.040.7886 0.8035 0.8353 0.8184 0.8065 0.7941 0.7994 0.7944
Llama-3.2-1B-Lit 51.26 56.1 0.7930 0.8279 0.8494 0.8296 0.8267 0.8216 0.8300 0.8314
Llama-3.2-1B-Tele-Lit 52.61 58.29 0.7978 0.8294 0.8519 0.8508 0.8316 0.8286 0.8318 0.8326
Llama-3.2-3B 39.29 25.5 0.8050 0.8207 0.8337 0.8282 0.7991 0.7904 0.8073 0.8040
Llama-3.2-3B-Lit 56.56 61.45 0.8049 0.8381 0.8486 0.8512 0.8331 0.8289 0.8291 0.8315
Llama-3.2-3B-Tele-Lit57.76 64.61 0.8097 0.8409 0.8479 0.8345 0.8309 0.8252 0.8314 0.8336
Llama-3-8B 51.75 55.67 0.8091 0.8171 0.8456 0.8240 0.8123 0.7998 0.8076 0.8066
Llama-3-8B-Lit 75.56 63.6 0.7903 0.8262 0.8548 0.8570 0.8343 0.8306 0.8332 0.8338
Llama-3-8B-Tele-Lit 78.56 66.98 0.7942 0.8286 0.8580 0.8605 0.8338 0.8317 0.8314 0.8330
GPT-4o 79.85 64.55 0.8648 0.8764 0.8544 0.8483 0.8259 0.8116 0.8723 0.8457
@@ VIII-C Instructions Fine-tuning
Training Settings. We maintain the training settings from Section VII with minor adjustments. Specifically, we decrease the batch size to 128k tokens, set the context length to 2048, and limit the number of epochs to 1. For this stage, we use two datasets: Alpaca and Open-Instruct [^11] https://huggingface.co/datasets/VMware/open-instruct. The combined datasets provide approximately 200k samples for training.
Results. In Table VI, we compare the Tele-Instruct versions of our models to their general instruct counterparts. Our instruction-tuned series outperforms the general instruct models, thereby showcasing their superior knowledge in telecom compared to their general-purpose counterparts.
VIII-D Benefits of Specialization
A central question we explore is how well domain specialization improves performance on downstream tasks beyond question answering. To investigate this, we focus on several literature-centric tasks: citation prediction, citation recommendation, title generation, abstract completion, and introduction-to-abstract conversion. Leveraging the Tele-Data dataset, we curate fine-tuning samples tailored to each of these tasks. We then fine-tune base models from the Llama family and compare their performance to the specialized Tele-LLMs, using a dedicated hold-out set excluded from training. Details on the fine-tuning process—including instruction datasets and hyperparameters—are provided in Appendix E.
For tasks involving text generation (e.g., citation sentence generation or abstract-to-title conversion), we adopted BERTScore-incorporating both precision and F1 measures- as our primary evaluation metric [38]. As shown in Table VIII-C, fine-tuned models (referred to by the suffix Lit) has the potential of outperforming GPT-4o on literature tasks (e.g., as is the case for LLama-3-8B) where domain-specific knowledge and graph structure are critical, while GPT-4o maintained an advantage in English language formulation. Notably, these fine-tuned models achieve competitive performance despite being significantly smaller than GPT-4o, demonstrating the efficiency benefits of targeted fine-tuning. Furthermore, the fine-tuned telecom models (referred to by the suffix Tele-Lit) consistently outperformed their generic counterparts across all tasks. This demonstrates how domain specialization fundamentally improves the model’s representations, aligning them more effectively with the telecom domain and enhancing performance on downstream tasks after task-specific fine-tuning. These results highlight the value of our developed Tele-LLMs series in enhancing performance on telecom-specific downstream tasks.
IX Conclusions
@@ Appendix D Qualitative Examples
Appendix E Literature Tasks Preparation
We create instruction datasets for literature tasks using our collected Tele-Data. The content of each paper generates instructions for title generation, abstract completion, and introduction-to-abstract tasks. Using citation information from Tele-Data, each edge creates a positive citation link prediction instruction. For balance, we replace the target paper in each edge with a random paper to create negative citation link prediction instructions, maintaining a 1:1 ratio. For every citation sentence associated with an edge, we develop an instruction that uses the titles and abstracts of both connected nodes to generate the citing sentence. For citation recommendation tasks, we randomly sample 10 negative nodes per edge to create a candidate set, requiring the model to identify the correct positive candidate.
During the fine-tuning process, we adopt the Q-LoRa approach [39], initializing the model in 8-bit precision and configuring the LoRa rank to 8. The LoRa layers are applied specifically to the $\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}$, and $\mathbf{O}$ matrices. To optimize training efficiency, we use a batch size of 8 along with a gradient accumulation step of 2. The learning rate is set to 0.0002, and optimization is performed using AdamW [40]. Furthermore, we assign a LoRa scaling hyperparameter $lora\_alpha$ of 32 and incorporate a dropout rate of 0.05 during fine-tuning.